# Exploring Frequency-Inspired Optimization in Transformer for Efficient Single Image Super-Resolution

### v3

@@ 6.4.6 Analysis of CRAFT’s Complexity
We compared CRAFT with SwinIR in terms of complexity using an input size of $128\times 128$, as shown in Table VI. The analysis considered parameters, FLOPs, GPU memory consumption, and average inference time. Compared to SwinIR, CRAFT has fewer parameters and FLOPs, and requires less memory consumption and inference time. Furthermore, we analyzed the complexity of our CRAFT framework and summarized the findings in Table VII. We observed that SRWAB contributes approximately 46% of the total complexity, while HFERB involves fewer convolution operations, resulting in reduced FLOPs. Moreover, the HFB module’s channel-wise attention effectively reduces the computational burden.
In addition, we investigated the impact of employing a mask mechanism in rectangle attention on the model’s efficiency, as illustrated in Table VIII. Our findings suggest that omitting the mask does not compromise performance but significantly boosts model inference speed. The results indicate that removing the mask leads to a reduction in inference time by approximately 37%, 41%, and 39% for magnification factors of 2, 3, and 4, respectively.
TABLE X: Efficacy of quantization methods across two stages, utilizing DBDC and PaC from PTQ4SR [27] as baselines to compare with our ADC and BR strategies. The symbol $\Delta$ indicates the improvement in PSNR compared with baseline methods at each stage. The dashed line separates the two stages for clarity.
@@
×$\times 4$ CRAFT-DBDC 23.69 22.18 21.90 20.24 22.49 21.36 20.77
CRAFT-ADC 26.21 (+2.52) 24.43 (+2.25) 24.05 (+2.15) 22.65 (+2.41) 23.23 (+0.74) 24.90 (+3.54) 23.21 (+2.44)
CRAFT-ADC 26.51 (+2.82) 24.61 (+2.43) 24.11 (+2.21) 22.77 (+2.53) 24.94 (+2.45) 25.06 (+3.70) 23.30 (+2.53)
@@
CRAFT-ADC with BR 28.95 (+0.01) 26.79 (+0.27) 26.27 (+0.24) 24.42 (+0.32) 27.07 (+1.44) 27.96 (+0.07) 24.87 (+0.11)
CRAFT-ADC with BR 29.46 (+0.52) 26.92 (+0.40) 26.32 (+0.29) 24.50 (+0.40) 27.03 (+1.40) 28.10 (+0.21) 24.97 (+0.21)
Figure 11: Clipping values with various clipping strategies applied to CRAFT with a magnification factor of 4. We visualize partial layers of the CRAFT to assess the effectiveness of our proposed ADC strategy. As the layer index increases, the values represent the results of deeper layers. From left to right, the results depict quantization at 8 bits, 6 bits, and 4 bits, respectively.
@@
Figure 12: Visualization of HFERB outputs across different quantization strategies: (a) Feature distribution of the full-precision model in spatial and frequency domains; (b) Output without FGO, showing significant frequency loss; (c) Output after applying FGO, with notably better frequency restoration.
6.5 Comparison with quantization strategies
TABLE XI: PSNR/SSIM comparison with transformer-based methods for the magnification factor of 4. All methods are quantized by the proposed general PTQ strategy. “W/A” represents weight and activation bit-width.
@@ 6.6.1 Effect of Frequency-guided Optimization
We conducted a detailed comparison with PTQ4SR [27], as shown in Table X, to highlight the benefits of our frequency-guided optimization (FGO). We used PTQ4SR [27] as a baseline for comparing our method’s effectiveness in dual clipping and boundary refinement. The strategies of DBDC and PaC, as described in PTQ4SR [27], were evaluated against our ADC and BR strategies. Initially, we compared our ADC strategy without FGO to PTQ4SR [27]’s DBDC and found that our ADC achieved better starting values, outperforming DBDC by up to 3.54dB. Following this, we assessed the performance of DBDC against our ADC with the FGO strategy included, clearly demonstrating that incorporating FGO into our strategy further improved performance, exceeding DBDC by up to 3.70dB.
Furthermore, we visualize the boundary values of the two different clipping strategies across 4-bit, 6-bit, and 8-bit quantizations with magnification factor of ×$\times 4$ in Fig. 11. We choose the output of CRFB, HFERB, SRWAB, and HFB as the probe. From the visual results, it’s evident that our ADC demonstrates a more constrained clipping range, providing a more precise initial value for subsequent training stages. This narrower range implies that a more accurate representation of the small range of values is sufficient to represent the full-precision counterparts.
@@ 6.6.2 General Transformer-based SR PTQ Strategy
Figure 13: Hardware performance of CRAFT and SwinIR on AMD Ryzen 5 5600U CPU using Lenovo XiaoXin Ari Plus+. The red diamond and blue triangle represent the speedup ratio of CRAFT and SwinIR. The quantized model consistently accelerates processing across various batch sizes.
@@ 6.6.3 Analysis of Quantized Acceleration
We have conducted a series of inference time evaluations on the CRAFT and SwinIR models using the Lenovo XiaoXin Air Plus+ device. This device is equipped with a 6-core AMD Ryzen 5 5600U CPU at 2.3GHz and 16 GB of RAM. For quantization, we used FBGEMM as the engine and AVX2 for the SIMD instruction set. Our tests involved using an input image size of 128x128 to perform 4x super-resolution. Due to the platform’s limited support capabilities, we restricted our quantization to 8 bits. The results are depicted in Figure 13. The term ”batch size” refers to the number of images processed simultaneously, and ”latency” indicates the processing time. We also highlighted the speedup using red diamonds and blue triangles. In terms of latency, the quantized model achieves lower latency compared to the full-precision model. Overall, across all experiments, the quantized model has demonstrated a promising speed improvement of up to 135%. Notably, compared to SwinIR, our CRAFT model consistently achieved greater speedup gains. However, it is important to note that the hardware platform supports only a limited number of quantization operators, which poses a challenge for transformer structures and limits further model acceleration.

### v4

@@ 6.4.6 Analysis of CRAFT’s Complexity
We compared CRAFT with SwinIR in terms of complexity using an input size of $128\times 128$, as shown in Table VI. The analysis considered parameters, FLOPs, GPU memory consumption, and average inference time. Compared to SwinIR, CRAFT has fewer parameters and FLOPs, and requires less memory consumption and inference time. Furthermore, we analyzed the complexity of our CRAFT framework and summarized the findings in Table VII. We observed that SRWAB contributes approximately 46% of the total complexity, while HFERB involves fewer convolution operations, resulting in reduced FLOPs. Moreover, the HFB module’s channel-wise attention effectively reduces the computational burden. In addition, we investigated the impact of employing a mask mechanism in rectangle attention on the model’s efficiency, as illustrated in Table VIII. Our findings suggest that omitting the mask does not compromise performance but significantly boosts model inference speed. The results indicate that removing the mask leads to a reduction in inference time by approximately 37%, 41%, and 39% for magnification factors of 2, 3, and 4, respectively.
6.5 Comparison with quantization strategies
TABLE X: Efficacy of quantization methods across two stages, using DBDC and PaC from PTQ4SR [26] as baselines for comparison with our ADC and BR strategies. The symbol $\Delta$ represents the PSNR improvement over the baseline methods at each stage. FGO denotes the proposed frequency-guided optimization. The dashed line separates the two stages for clarity.
@@ 6.5 Comparison with quantization strategies
×$\times 4$ DBDC [26] 23.69 22.18 21.90 20.24 22.49 21.36 20.77
ADC (Ours) 26.21 (+2.52) 24.43 (+2.25) 24.05 (+2.15) 22.65 (+2.41) 23.23 (+0.74) 24.90 (+3.54) 23.21 (+2.44)
ADC (Ours) 26.51 (+2.82) 24.61 (+2.43) 24.11 (+2.21) 22.77 (+2.53) 24.94 (+2.45) 25.06 (+3.70) 23.30 (+2.53)
@@
BR (Ours) 28.95 (+0.01) 26.79 (+0.27) 26.27 (+0.24) 24.42 (+0.32) 27.07 (+1.44) 27.96 (+0.07) 24.87 (+0.11)
BR (Ours) 29.46 (+0.52) 26.92 (+0.40) 26.32 (+0.29) 24.50 (+0.40) 27.03 (+1.40) 28.10 (+0.21) 24.97 (+0.21)
Figure 12: Clipping strategies comparison between PTQ4SR [26] and our method applied to CRAFT with a magnification factor of 4. The layer index represents progressively deeper layers in the model. From left to right, the results show quantization at 8 bits, 6 bits, and 4 bits, respectively.
@@
Figure 13: Visualization of HFERB outputs under different quantization strategies. FGO refers to the proposed frequency-guided optimization. (a) Feature distribution of the full-precision model in both spatial and frequency domains. (b) Output without FGO, showing significant frequency loss. (c) Output after applying FGO, demonstrating notably improved frequency restoration.
TABLE XI: PSNR/SSIM comparison with transformer-based methods for the magnification factor of 4. All methods are quantized by the proposed general PTQ strategy. “W/A” represents weight and activation bit-width, while FGO refers to the proposed frequency-guided optimization.
@@ 6.6.1 Effect of Frequency-guided Optimization
We conducted a detailed comparison with PTQ4SR [26], as shown in Table X, to highlight the benefits of our frequency-guided optimization (FGO). We used PTQ4SR [26] as a baseline for comparing our method’s effectiveness in dual clipping and boundary refinement. The strategies of DBDC and PaC, as described in PTQ4SR [26], were evaluated against our ADC and BR strategies. Initially, we compared our ADC strategy without FGO to PTQ4SR [26]’s DBDC and found that our ADC achieved better starting values, outperforming DBDC by up to 3.54dB. Following this, we assessed the performance of DBDC against our ADC with the FGO strategy included, clearly demonstrating that incorporating FGO into our strategy further improved performance, exceeding DBDC by up to 3.70dB. Furthermore, we visualize the boundary values of the two different clipping strategies across 4-bit, 6-bit, and 8-bit quantizations with magnification factor of 4 in Fig. 12. We choose the output of CRFB, HFERB, SRWAB, and HFB as the probe. From the visual results, it’s evident that our ADC demonstrates a more constrained clipping range, providing a more precise initial value for subsequent training stages. This narrower range implies that a more accurate representation of the small range of values is sufficient to represent the full-precision counterparts.
@@ 6.6.2 General Transformer-based SR PTQ Strategy
Figure 14: Hardware performance of CRAFT and SwinIR on AMD Ryzen 5 5600U CPU. The red diamond and blue triangle represent the speedup ratio of CRAFT and SwinIR. The quantized model consistently accelerates processing across various batch sizes.
TABLE XII: Comparison of PTQ with different training strategies. “Regular” refers to full-precision training, while “AMP” stands for Automatic Mixed Precision training. “Prec.” denotes the precision value.
Model Prec. Model Size Set14 BSD100 Manga109
Regular FP32 3.37M 28.85/.7872 27.72/.7418 31.18/.9168
AMP FP16 1.81M 28.83/.7871 27.71/.7416 31.13/.9160
FP8 1.51M 28.15/.7708 27.29/.7266 29.16/.8888
PTQ INT8 1.13M 28.80/.7850 27.68/.7397 31.06/.9141
@@ 6.6.3 Analysis of Quantized Acceleration
The primary goal of this evaluation was to assess the effects of quantization on resource-constrained platforms, such as portable PC CPUs, which more closely resemble the typical deployment scenarios for quantized models. While 8-bit quantization is also applicable to GPUs, our initial focus was to investigate the performance of quantized models on hardware with more limited resources. Due to limitations in current hardware, which offers limited support for fully quantized models, especially for lower-bit operations (e.g., 6-bit or 4-bit), we restricted our quantization to 8-bit for this study. The tests were performed on input images sized $128\times 128$, applying a ×$\times 4$ super-resolution. The acceleration results are presented in Fig. 14, where we compared our CRAFT model with SwinIR [3] to evaluate the effectiveness of our quantization technique. The quantized INT8 model consistently exhibited lower latency compared to the full-precision FP32 model, achieving a speed improvement of up to 24%. However, the actual speedup fell short of theoretical expectations. We believe this is due to several factors, including limitations in INT8 batch matrix multiplication and the presence of non-quantizable operations, such as softmax and LayerNorm, which reduce the overall acceleration potential of INT8 quantization during inference. To further highlight the advantages of PTQ, we performed additional comparisons with models trained using the Automatic Mixed Precision (AMP) strategy, which is widely used in model training [57, 33, 58]. As shown in Table XII, the models were trained with AMP in both FP16 and FP8111FP8 training was conducted using NVIDIA’s Transformer Engine, which only supports converting the linear layers to FP8, while other layers remain in FP16.. While AMP-trained FP16 models demonstrated only a slight performance decrease compared to full-precision models, FP8 showed significant degradation. In contrast, PTQ achieved notably smaller model sizes while maintaining competitive performance. These results emphasize PTQ’s superior efficiency, offering clear advantages in terms of both performance and model size over floating-point models.